can you please tell me whetther these are true or false and how you worked them out?
1. 3(x+2)=3x+5
2. 5(x-4)=5X-20
2. 2(4x+3=14x

Answer by stanbon(75887)

can you please tell me whetther these are true or false and how you worked them out?
1. 3(x+2)=3x+5
False:
3(x+2) = 3*x + 3*2 = 3x+6
---------------------------
2. 5(x-4)=5X-20
True:
---------------------------
2. 2(4x+3)=14x
False:
2(4x+3) = 2*4x + 2*3 = 8x+6
===================
Cheers,
Stan H.
===========